A net force of 126N (S) is applied to a 70 Kg sprinter. Determine the acceleration of the sprinter


1
Expert's answer
2020-11-06T10:12:31-0500

We can find the acceleration of the sprinter from the Newton's Second Law of Motion:


"F_{net}=ma,""a=\dfrac{F_{net}}{m}=\dfrac{126\ N\ [S]}{70\ kg}=1.8\ \dfrac{m}{s^2}\ [S]."

Answer:

"a=1.8\ \dfrac{m}{s^2}\ [S]."